# Metric Differential Privacy at the User-Level

### v1

@@ Appendix D Omitted Proofs from Section 6
D.1. Proof of Lemma 6.2
See 6.2 As the sensitivity of the query is bounded by $\|F\|_{2}$, is easy to show (e.g. (Dwork et al., 2014)) that adding $d$-dimensional Gaussian noise with width $\|F\|_{2}\frac{r\sqrt{1.25\ln\frac{1}{\delta}}}{\alpha}$ in each coordinate will satisfy $(\frac{\alpha}{r},\delta)$ local $d_{\textsf{EM}}$-DP. The standard deviation in each coordinate of $\hat{q}$ is thus $\|F\|_{2}\frac{r\sqrt{1.25\ln\frac{1}{\delta}}}{\alpha\sqrt{n}}$, and this gives the desired expected error.
D.2. Proof of Theorem 6.5
@@
(25) $\leq r\sqrt{st\mathbb{E}[\|\tilde{F}-\tilde{K}\|_{2}^{2}]}+\sqrt{s\mathbb{E}[\|(\tilde{F}-\tilde{K})P\|_{2}^{2}]}.$

### v2

@@ Appendix D Omitted Proofs from Section 6
D.1 Proof of Lemma 6.2
See 6.2 As the sensitivity of the query is bounded by $\|F\|_{2}$, is easy to show (e.g. Dwork et al. (2014)) that adding $d$-dimensional Gaussian noise with width $\|F\|_{2}\frac{r\sqrt{1.25\ln\frac{1}{\delta}}}{\alpha}$ in each coordinate will satisfy $(\frac{\alpha}{r},\delta)$ local $d_{\textsf{EM}}$-DP. The standard deviation in each coordinate of $\hat{q}$ is thus $\|F\|_{2}\frac{r\sqrt{1.25\ln\frac{1}{\delta}}}{\alpha\sqrt{n}}$, and this gives the desired expected error.
D.2 Proof of Theorem 6.5
@@
$\leq r\sqrt{st\mathbb{E}[\|\tilde{F}-\tilde{K}\|_{2}^{2}]}+\sqrt{s\mathbb{E}[\|(\tilde{F}-\tilde{K})P\|_{2}^{2}]}.$ (26)